sqrt(3x+4) - sqrt (2x+4) =2
Answer by stanbon(75887)

sqrt(3x+4) - sqrt (2x+4) =2
---
sqrt(3x+4) = sqrt(2x+4) + 2
-----
Square both sides to get:
3x+4 = 2x+4 + 4*sqrt(2x+4) + 4
x - 4 = 4sqrt(2x+4)
----
Square both sides:
x^2 - 8x + 16 = 16(2x+4)
----
x^2 - 40x - 48 = 0
----
x = [40 +- sqrt(1600-4*-48)]/2
----
x = [40 +- sqrt(1792)]/2
----
x = [20 +- 42.33]/2
x = 61.33/2 or x = -22.33/2
Positive solution:
x = 30.67
-------------------
Check to see if that solution is extraneous.
-------------------
Cheers,
Stan H.